# photon/.style=decorate, decoration=snake, draw=black

### v1

@@ IX.2 CFF Extraction Results
Table 5: The breakdown of the QDNN error contributions are listed for model 0-3, as described in the text, for each CFF and DVCS term.
IX.3 Addressing QDNN Algorithmic Error
To systematically investigate the sources of algorithmic error in the QDNN, we implemented three new QDNN models with targeted modifications. Model 0 is our baseline QDNN, which we used for all previous CFF extraction. Even though it isn’t the most optimal, it provides a simple quantum model to demonstrate the inherent quantum advantage over CDNNs in CFF extraction. Therefore, since it exhibits higher algorithmic error than our CDNN model, it is a useful starting point for identifying areas of improvement. To address these issues, we introduce Models 1-3, each focusing on a specific refinement to improve stability and reduce algorithmic error. Model 1 modifies the strongly entangling layers by adding an entanglement range parameter, limiting interactions to nearest neighbors in some cases rather than fully entangling all qubits at every layer. This aims to balance expressivity and stability, reducing excessive parameter correlations that can destabilize optimization. Model 2 focuses on addressing the barren plateau problem, where gradients vanish exponentially as the circuit depth increases. To mitigate this, Model 2 employs layerwise training, where the network begins with a shallow quantum circuit and progressively increases the number of layers over training. Additionally, it introduces depth-scaled initialization, ensuring that each layer’s parameter magnitudes decrease as depth increases. This prevents excessive parameter magnitudes in deeper layers while maintaining trainable gradients throughout optimization. Model 3 refines the small-angle initialization, which is already known to help prevent barren plateaus at initialization but may overly restrict the search space, leading to increased algorithmic error. Instead of reducing the initial weight magnitudes across layers, Model 3 rescales small-angle initialization to maintain better trainability without completely restricting parameter exploration.
To test these new models, we perform the CFF extraction on 1000 identical noisy replicas of Set 144 (see Figs. 7 and 9) using each of these four models and display the algorithmic error results in Table 5. Comparing the results of Model 0 with Figure 11, we observe that the algorithmic errors for this kinematic bin align relatively well with the average algorithmic errors of the extracted CFFs across all bins. Performing the extraction with the improved QDNN models, the algorithmic error improves significantly for all three models. Table 5 indicates that Model 1 produces the lowest algorithmic errors for $\Re e\mathcal{H}$, $\Re e\mathcal{E}$, and $\Re e\widetilde{\mathcal{H}}$, while Model 2 produces the lowest algorithmic error for $DVCS$. Although these new models aim to improve QDNN algorithmic error, we also examine their effect on the other three error metrics (methodological error, precision, and accuracy) and display the results in Table 5. As expected, the improved algorithmic error results in better precision across Models 1-3, with the improved precision following a similar pattern to the algorithmic error. We find that these new models do not significantly affect the methodological error or accuracy.
The QDNN algorithmic error improvements observed from Models 1-3 are more comparable to the algorithmic errors from the CDNN extraction (see Fig. 11) and suggest key strategies for reducing algorithmic error and improving precision in quantum deep learning models for CFF extraction. Specifically, by refining entanglement structure, implementing layerwise training, and adjusting initialization schemes, we demonstrate that targeted modifications can significantly enhance training stability and overall model reliability. These tests provide a foundation for future studies to develop more optimized QDNN architectures, particularly by exploring adaptive entanglement strategies, hybrid classical-quantum initialization techniques, and quantum-specific optimizers.
X Conclusions
In this work, we explored the application of QDNNs for the extraction of CFFs from Deeply Virtual Compton Scattering (DVCS) pseudodata and performed a systematic comparison with CDNNs under matched model complexity. To enable a direct head-to-head comparison of algorithmic performance, we deliberately constrained the QDNN architecture, ensuring parity in complexity and hyperparameter tuning between the quantum and classical models. This allowed us to isolate the effect of quantum computational mechanisms—such as entanglement and superposition—from differences in network capacity or training depth.
We developed a novel set of quantum qualifiers to identify the specific regimes—characterized by data dimensionality, function complexity, and training sparsity—where QDNNs offer the greatest advantage. These metrics extend beyond traditional performance evaluation by predicting scenarios where QDNNs are expected to outperform based on intrinsic data complexity. Through both classification and regression tasks, these quantum metrics consistently revealed enhanced performance of QDNNs in learning nontrivial correlations and high-dimensional feature spaces more efficiently than their classical counterparts.
Our application of QDNNs to DVCS pseudodata demonstrated significantly improved extraction of CFFs, with QDNNs achieving higher accuracy, greater precision, and lower total methodological error compared to CDNNs. Notably, while total error was reduced, QDNNs exhibited larger algorithmic error in some cases, likely arising from quantum-specific challenges such as restricted initialization schemes, gradient propagation limitations, and entanglement-induced parameter correlations. These findings highlight areas for targeted optimization in future QDNN architectures.
To maintain parity for comparison, the QDNNs in this work were not fully optimized—nor was a QDNN-specific extraction scheme developed. As such, the true potential of quantum-enhanced inference remains only partially tapped. Future work should focus on building tailored QDNN extraction frameworks, exploring alternative initialization methods such as amplitude embedding, incorporating quantum-specific optimizers (e.g., quantum natural gradient descent), and systematically investigating how entanglement influences learning dynamics and generalization.
Ultimately, this study provides a first step toward quantum-enhanced parameter extraction in hadronic physics. The demonstrated advantage of QDNNs in capturing high-dimensional quantum correlations positions them as a promising tool for future analyses—not only in DVCS but across a range of deep exclusive processes. As quantum hardware continues to mature, and as hybrid quantum-classical learning algorithms improve in stability and scalability, QDNNs may become essential in realizing high-precision, model-independent extractions of hadronic observables. Future work will extend these methods to real experimental data, further developing quantum deep learning as a transformative approach for unraveling the structure of matter at the most fundamental level.
XI Acknowledgment
Appendix A Tables
(26) (27) (28) (29) (30) (31)
$\mathfrak{N}$ 0.982 0.995 0.996 0.939 0.775 0.983
$\Phi$ 26 29 36 27 25 28
$\mathfrak{D}$ 1.637 1.491 1.679 1.488 1.525 1.531
$\mathfrak{M}$ 0.648 0.762 0.504 0.833 0.784 0.622
(26) (27) (28) (29) (30) (31)
$\mathfrak{N}$ 0.986 0.994 0.994 0.942 0.778 0.983
$\Phi$ 31 31 36 28 25 30
$\mathfrak{D}$ 1.618 1.533 1.684 1.518 1.547 1.553
$\mathfrak{M}$ 0.667 0.607 0.453 0.752 0.791 0.625
Table 7: Function characteristic data for regression target functions [Eqs. (26)-(31)] with $0.25\sigma$ noise level.
(26) (27) (28) (29) (30) (31)
$\mathfrak{N}$ 0.997 0.997 1.000 0.963 0.879 0.977
$\Phi$ 33 35 36 29 33 26
$\mathfrak{D}$ 1.609 1.638 1.715 1.517 1.586 1.665
$\mathfrak{M}$ 0.468 0.458 0.497 0.541 0.585 0.485
Table 8: Function characteristic data for regression target functions [Eqs. (26)-(31)] with $1\sigma$ noise level.
Epochs (26) (27) (28) (29) (30) (31)
10 3.818 5.373 5.929 6.164 6.921 3.857
50 2.348 4.272 5.818 4.592 4.688 3.166
100 1.700 3.670 5.524 3.183 3.841 2.980
250 0.249 1.790 4.784 2.286 3.277 1.366
500 0.382 1.162 4.345 1.879 3.295 0.922
Epochs (26) (27) (28) (29) (30) (31)
10 3.151 4.603 5.877 4.490 4.350 3.416
50 0.183 2.859 5.790 3.310 3.929 2.183
100 0.191 2.335 5.682 1.833 3.604 1.616
250 0.204 2.165 5.576 1.291 3.341 1.589
500 0.209 0.975 5.462 1.171 3.289 0.433
Epochs (26) (27) (28) (29) (30) (31)
10 3.824 5.373 5.928 6.164 6.921 3.859
50 2.259 4.242 5.772 4.624 4.705 3.137
100 1.937 3.689 5.531 3.065 3.868 3.007
250 0.554 2.754 5.029 2.244 3.410 1.666
500 0.622 1.572 4.686 2.077 3.441 1.188
Epochs (26) (27) (28) (29) (30) (31)
10 3.697 4.620 5.964 4.503 4.301 3.438
50 0.388 2.978 5.748 3.337 4.064 3.030
100 0.444 2.382 5.698 2.030 3.800 1.833
250 0.445 1.642 5.613 1.658 3.518 1.081
500 0.581 1.400 5.493 1.455 3.582 0.946
Epochs (26) (27) (28) (29) (30) (31)
10 3.865 5.371 5.923 6.164 6.922 3.881
50 3.431 4.228 5.711 4.940 4.905 3.663
100 2.593 3.500 5.602 3.707 4.515 3.459
250 2.633 3.842 5.855 3.596 4.633 2.985
500 2.215 3.390 5.528 3.015 4.429 2.904
Epochs (26) (27) (28) (29) (30) (31)
10 3.697 4.815 5.913 4.785 5.008 3.633
50 1.453 3.352 5.905 3.906 4.611 3.508
100 1.661 3.069 5.781 3.430 4.480 3.270
250 2.030 2.907 5.778 3.007 4.449 2.464
500 2.635 3.341 4.832 3.222 4.471 3.052
0.75 to 1 25 to 35 1.45 to 1.75 0.45 to 0.85 5000 to 35000
Epochs $\mathfrak{N}$ $\Phi$ $\mathfrak{D}$ $\mathfrak{M}$ $\mathfrak{F}$
10 -2.43 -0.0386 -1.69 1.12 7.06E-06
50 0.209 -0.0257 -1.89 0.725 -8.55E-06
100 1.36 -0.0327 -2.66 0.87 -3.04E-05
250 0.715 -0.0115 -1.26 0.379 -5.59E-06
500 0.746 -0.0476 -1.64 1.1 -2.26E-05
Average 0.12 -0.0312 -1.83 0.839 -1.20E-05
Average $\Delta\Xi$ over domain 0.03 -0.312 -0.549 0.337 -0.36
Table 15: Slopes of linear regressions on the quantum outperformance $\Xi$ against the function characteristics [nonlinearity ($\mathfrak{N}$), frequency complexity ($\Phi$), fractal dimension ($\mathfrak{D}$), mutual information ($\mathfrak{M}$), and Fourier transform complexity ($\mathfrak{F}$)] for DNN target function regression. The approximate range in the function characteristics, the average slopes across the different epochs, and the average change in $\Xi$ over the function characteristic domain are reported.
Epochs $\mathfrak{N}$ $\Phi$ $\mathfrak{D}$ $\mathfrak{M}$ $\mathfrak{F}$
10 0.829 0.573 0.384 0.524 0.079
50 0.008 0.306 0.596 0.267 0.12
100 0.104 0.155 0.368 0.13 0.467
250 0.032 0.023 0.096 0.027 0.022
500 0.022 0.246 0.102 0.144 0.228
Average 0.199 0.261 0.309 0.218 0.183
Table 16: $R^{2}$ values of linear regressions on the quantum outperformance $\Xi$ against the function characteristics [nonlinearity ($\mathfrak{N}$), frequency complexity ($\Phi$), fractal dimension ($\mathfrak{D}$), mutual information ($\mathfrak{M}$), and Fourier transform complexity ($\mathfrak{F}$)] for DNN target function regression. The average $R^{2}$ values across the different epochs are reported.

### v2

@@ A.3 Tables
Full 0.127 0.161 0.175 0.000502 0.0408 0.0470 0.0356 0.000157 0.402 0.578 0.864 0.00511 0.773 0.0327 0.00018 0.00264
Table 19: Results from the optimized CFF extraction from the real experimental data subset performed in Sec. VII. For each set of kinematics, the mean and standard deviation of the FQDNN-predicted CFFs are listed.
Set $k$ $Q^{2}$ $x_{B}$ $t$ Mean $\Re e\mathcal{H}$ Std $\Re e\mathcal{H}$ Mean $\Re e\mathcal{E}$ Std $\Re e\mathcal{E}$ Mean $\Re e\widetilde{\mathcal{H}}$ Std $\Re e\widetilde{\mathcal{H}}$ Mean $DVCS$ Std $DVCS$
108 5.750 1.790 0.244 -0.110 -1.304 0.714 -1.604 0.798 -1.548 0.848 -0.0137 0.0200
109 5.750 1.790 0.244 -0.150 -0.222 0.498 -0.540 0.825 0.058 0.757 0.0201 0.0113
110 5.750 1.790 0.244 -0.200 -0.934 0.503 -0.962 0.544 -1.448 0.726 0.0080 0.0085
111 5.750 1.790 0.244 -0.260 -0.577 0.543 -0.724 0.635 -0.208 1.085 0.0196 0.0076
112 5.750 1.800 0.244 -0.340 -0.465 0.905 -0.731 0.912 0.160 0.836 0.0222 0.0080
113 5.750 1.800 0.244 -0.450 -0.508 1.640 -0.400 1.404 -0.240 0.596 0.0166 0.0090
114 5.750 2.120 0.244 -0.110 1.270 0.826 1.376 0.855 1.405 0.901 -0.0234 0.0170
115 5.750 2.120 0.245 -0.150 2.950 0.990 2.909 0.908 2.937 1.127 0.0406 0.0136
116 5.750 2.120 0.245 -0.200 4.590 1.532 3.880 1.371 3.496 1.437 0.0453 0.0176
117 5.750 2.120 0.245 -0.260 5.603 1.775 4.825 1.587 3.557 1.468 0.0573 0.0170
118 5.750 2.120 0.245 -0.340 9.494 3.111 8.464 2.769 5.397 2.212 0.0723 0.0223
119 5.750 2.120 0.245 -0.450 0.538 1.557 0.588 1.438 0.119 0.931 0.0193 0.0101
120 5.750 1.940 0.274 -0.110 -0.594 1.161 -0.973 1.988 -1.083 1.833 0.0261 0.0280
121 5.750 1.940 0.274 -0.150 -1.234 0.626 -1.707 0.754 -1.894 0.795 0.0022 0.0119
122 5.750 1.940 0.274 -0.200 -0.700 0.439 -1.177 0.575 -0.523 0.674 0.0028 0.0078
123 5.750 1.940 0.274 -0.260 -1.132 0.460 -1.286 0.491 -1.729 0.570 0.0091 0.0069
124 5.750 1.940 0.275 -0.340 -0.152 0.504 -0.218 0.573 0.392 1.135 0.0086 0.0052
125 5.750 1.940 0.275 -0.450 -0.159 0.756 -0.129 0.919 -0.305 0.613 0.0133 0.0040
126 5.750 2.350 0.274 -0.110 0.911 1.706 1.269 2.371 1.203 2.245 -0.0420 0.0366
127 5.750 2.350 0.275 -0.150 0.580 0.911 0.701 1.094 0.718 1.204 0.0038 0.0105
128 5.750 2.350 0.275 -0.200 1.874 0.899 1.975 0.874 2.159 1.023 0.0150 0.0073
129 5.750 2.350 0.275 -0.260 3.161 1.189 3.095 1.138 2.886 1.323 0.0288 0.0077
130 5.750 2.350 0.275 -0.340 3.104 1.320 2.424 1.122 1.522 0.858 0.0246 0.0067
131 5.750 2.350 0.275 -0.450 7.237 2.637 6.298 2.351 3.682 1.580 0.0340 0.0102
133 5.750 2.100 0.304 -0.150 -2.007 1.511 -3.181 2.303 -3.098 2.231 0.0263 0.0319
134 5.750 2.100 0.304 -0.200 -0.690 0.453 -1.249 0.698 -1.231 0.619 0.0040 0.0062
135 5.750 2.100 0.304 -0.260 -0.651 0.437 -0.938 0.567 -1.174 0.719 0.0058 0.0059
136 5.750 2.100 0.304 -0.340 -0.662 0.322 -0.802 0.368 -1.464 0.499 0.0053 0.0040
137 5.750 2.100 0.305 -0.450 -0.329 0.390 -0.368 0.424 -0.547 0.845 0.0116 0.0029
139 5.750 2.580 0.304 -0.150 -0.852 2.441 -1.130 3.231 -1.087 3.148 0.0128 0.0421
140 5.750 2.580 0.304 -0.200 0.094 0.545 0.143 0.847 0.085 0.918 -0.0003 0.0049
141 5.750 2.580 0.305 -0.260 2.103 0.771 2.269 0.740 2.566 0.799 0.0100 0.0039
142 5.750 2.580 0.305 -0.340 1.524 0.932 1.620 0.909 1.085 0.907 0.0106 0.0038
143 5.750 2.580 0.305 -0.450 2.056 1.553 1.510 1.189 0.790 0.741 0.0135 0.0042
145 5.750 2.230 0.335 -0.200 -0.086 0.381 -0.197 0.933 -0.271 1.036 -0.0060 0.0048
146 5.750 2.230 0.335 -0.260 -0.658 0.456 -1.031 0.669 -1.619 0.773 0.0080 0.0059
147 5.750 2.230 0.335 -0.340 -0.429 0.393 -0.548 0.484 -0.867 0.780 0.0086 0.0040
148 5.750 2.230 0.335 -0.450 -1.064 0.311 -1.210 0.338 -2.108 0.395 0.0044 0.0026
150 5.750 2.780 0.335 -0.200 -0.914 0.635 -1.603 0.866 -1.505 0.888 0.0026 0.0054
151 5.750 2.780 0.335 -0.260 -0.052 0.385 -0.081 0.768 -0.223 0.864 0.0015 0.0027
152 5.750 2.780 0.335 -0.340 1.135 0.615 1.442 0.658 1.700 0.760 0.0075 0.0022
153 5.750 2.780 0.335 -0.450 0.784 0.845 0.989 0.995 0.429 0.650 0.0060 0.0020
154 5.750 2.350 0.365 -0.200 -1.929 1.139 -5.204 2.382 -4.058 2.066 0.0097 0.0110
155 5.750 2.350 0.365 -0.260 -0.902 0.543 -1.813 0.936 -1.961 0.998 0.0047 0.0057
156 5.750 2.350 0.365 -0.340 -0.340 0.412 -0.507 0.591 -0.836 0.943 0.0061 0.0040
157 5.750 2.350 0.365 -0.450 -0.373 0.284 -0.462 0.343 -0.983 0.606 0.0036 0.0027
158 5.750 2.970 0.364 -0.200 -0.330 0.875 -0.857 2.353 -0.649 1.728 -0.0014 0.0055
159 5.750 2.970 0.365 -0.260 0.007 0.331 -0.015 0.970 0.083 1.067 0.0037 0.0026
160 5.750 2.970 0.365 -0.340 0.583 0.400 1.137 0.603 1.293 0.707 0.0021 0.0017
161 5.750 2.970 0.365 -0.450 0.976 0.391 1.283 0.444 1.943 0.444 0.0036 0.0011
163 5.750 2.480 0.400 -0.260 0.140 0.521 0.375 1.376 0.293 1.123 -0.0001 0.0047
164 5.750 2.480 0.399 -0.340 -0.193 0.524 -0.323 0.821 -0.434 1.061 0.0057 0.0045
165 5.750 2.480 0.399 -0.450 -0.699 0.312 -0.905 0.385 -1.481 0.574 0.0018 0.0024
166 5.750 3.120 0.391 -0.210 -1.311 1.796 -3.921 4.696 -2.405 3.127 0.0191 0.0134
167 5.750 3.180 0.399 -0.260 -1.064 0.742 -2.757 1.586 -2.147 1.324 0.0002 0.0038
168 5.750 3.180 0.400 -0.340 -0.347 0.273 -1.059 0.645 -1.434 0.731 0.0023 0.0015
169 5.750 3.180 0.400 -0.450 0.073 0.139 0.155 0.269 0.189 0.795 0.0032 0.0009
171 5.750 2.960 0.449 -0.340 0.248 0.424 0.757 1.164 0.549 0.867 0.0010 0.0027
172 5.750 3.050 0.461 -0.450 -0.114 0.186 -0.276 0.400 -0.411 0.521 0.0016 0.0009
173 5.750 3.500 0.430 -0.280 -1.382 1.844 -3.202 3.750 -2.327 2.896 -0.0006 0.0098
174 5.750 3.630 0.451 -0.340 -0.072 0.345 -0.256 1.167 -0.168 0.723 0.0006 0.0016
175 5.750 3.770 0.475 -0.450 0.108 0.174 0.279 0.440 0.275 0.405 0.0010 0.0006